# RefineBridge: Generative Bridge Models improve financial forecasting by Foundation Models

### v1

@@ 1 Introduction
Inspired by the scaling laws observed in foundation models for natural language processing [grattafiori2024llama] and computer vision [dosovitskiy2020image], Time Series Foundation Models (TSFM) [chronos, moirai, timemoe, rasul2024lagllama, liutimer] have extended this scaling paradigm to time series forecasting, with models containing millions to billions of parameters. By scaling the transformer architecture [vaswani2017attention] and pre-training on diverse time series datasets, TSFMs can learn universal temporal representations and have achieved state-of-the-art performance across a wide range of forecasting tasks.
@@
Despite the presence of financial time series in the training data of TSFMs, their prediction remains particularly challenging due to domain-specific statistical properties [fintsb2025benchmark]. In particular, financial data often exhibit near-zero autocorrelation, strong non-stationarity, and heteroskedasticity which shifts across market regimes. As a result, TSFMs tend to default to mean-reverting forecasts and output the average of the context window irrespective of market conditions. This leads to predictions that miss higher-frequency variations, which frequently carry tradable signals.
@@
Fig. 1: RefineBridge learns an optimal transport map from TSFM predictions to the ground truth via Schrödinger Bridge. At inference time, RefineBridge progressively refines the TSFM forecasts along sampling steps.
@@
Low-Rank Adaptation (LoRA) [lora] has emerged as a popular parameter-efficient method for adapting TSFMs to downstream data domains [gupta2024beyond]. LoRA-based adapters introduce low-rank update matrices with significantly fewer trainable parameters than the original foundation model, enabling parameter-efficient fine-tuning. However, when applied to financial data, LoRA is prone to overfitting and can even degrade TSFM performance (see Table 1). Specifically, LoRA typically adopts the same network architecture and objective function as the original model, rather than introducing a complementary objective to enhance predictions, inherently limiting its performance.
@@
To address these issues, we propose RefineBridge, a novel paradigm for improving the performance of TSFMs on financial time series forecasting. In contrast to LoRA, which functions as a fine-tuning technique, RefineBridge operates as an independent post-processing module. Specifically, it treats TSFM forecasts as the prior distribution and iteratively refines them through Schrödinger Bridge optimal transport [bridgetts, wang2025framebridge], conditioned on the context window. As a probabilistic generative model, RefineBridge introduces a complementary training objective which captures the optimal stochastic transport maps from the prior to the target distribution. By using the coarse predictions by TSFMs as a prior, RefineBridge learns to model fine-grained details, which are crucial in financial time series forecasting. During the sampling stage, RefineBridge generates a refinement trajectory which aligns TSFM forecasts with the unique statistical properties of downstream financial data, all while keeping the original TSFM parameters unchanged, as illustrated in Fig. 1. The main contributions of this work are as follows.
@@ 2 Related Work
Time Series Foundation Models. Recent advancements in deep learning have led to the development of various TSFMs [chronos, moirai, timemoe, rasul2024lagllama, liutimer], with Chronos [chronos], Moirai [moirai], and Time-MoE [timemoe] representing three distinct architectures with different training objective functions. Chronos [chronos] adapts the T5 [t5] architecture by discretizing continuous time series into tokens and treating forecasting as a sequence-to-sequence task, where future tokens are generated auto-regressively. Moirai [moirai] instead uses mixture distributions and a masked encoder architecture with continuous embeddings. At a larger scale, Time-MoE [timemoe] is the first TSFM to reach the billion-parameter level through a sparse mixture-of-experts architecture.
Low-rank Adaptation. Full fine-tuning of TSFMs for specific downstream datasets is computationally expensive due to their large parameter count. To address this, LoRA-based adapters [lora, kopiczko2024vera, gu2025tera] introduce low-rank weight update matrices to greatly reduce the number of trainable parameters during fine-tuning. However, in practice, LoRA struggles when applied to financial data, as shown in Table 1.
@@ 3 RefineBridge
Unlike diffusion models which map noise to data, Schrödinger Bridge learns data-to-data transport maps by computing the most probable stochastic evolution between two probability distributions under a reference diffusion process [sb_fbsde, diffusion_schrodinger_bridge]. This makes SB well-suited for refining TSFM predictions, which can be viewed as neural priors learned by the internal representation of TSFMs, as illustrated in Fig. 1. Even though TSFMs often struggle with financial time series, their forecasts still contain a coarse estimate (the neural prior) of the prediction target. As highlighted by [wang2025framebridge], compared to noise-to-data diffusion models, SB-based models can capture the target distribution from a prior distribution, making it possible to build post-processing refinement models for pretrained foundation models.
@@ 3.1 Bridging TSFM Predictions and Ground Truth
Following prior work [bridgetts, wang2025framebridge, li2025bridge], we adopt a tractable solution for paired data $(\mathbf{x}_{0},\mathbf{x}_{T})$. We set the reference stochastic differential equation (SDE) to have zero drift ($f(t)=0$) and a time-varying diffusion defined as
$d\mathbf{x}_{t}=g(t)d\mathbf{w}_{t}$, where $g^{2}(t)=\beta_{0}+t(\beta_{1}-\beta_{0})$. Here, $f:\mathbb{R}^{H}\times[0,T]\to\mathbb{R}^{H}$ is the vector-valued drift term, $g:[0,T]\to\mathbb{R}$ is scalar-valued diffusion coefficient, and $\mathbf{w}_{t}\in\mathbb{R}^{H}$ denotes a standard Wiener process. This SDE defines the continuous-time diffusion process upon which score-based generative models (SGMs) [sgm] are built.
In the general SB framework with linear drift, the scaling coefficients are given by $\alpha_{t}=e^{\int_{0}^{t}f(\tau)d\tau}$ and $\bar{\alpha}_{t}=e^{-\int_{t}^{1}f(\tau)d\tau}$. We employ the Bridge-gmax scheduler proposed in [bridgetts], to simplify these to $\alpha_{t}=\bar{\alpha}_{t}=1$ under $f(t)=0$. The corresponding variance schedules are $\sigma_{t}^{2}=\int_{0}^{t}g^{2}(\tau)d\tau=\beta_{0}t+\frac{1}{2}(\beta_{1}-\beta_{0})t^{2}$ and $\bar{\sigma}_{t}^{2}=\sigma_{1}^{2}-\sigma_{t}^{2}.$ The resulting bridge marginal distribution is Gaussian
@@ 3.2 Complementary Training Objective
When LoRA [lora] is applied to a TSFM for downstream adaptation, it typically retains the training objective of the base model, thereby potentially inheriting its limitations. This includes maximizing token-level log-likelihood in Chronos [chronos], optimizing a mixture distribution log-likelihood in Moirai [moirai], and minimizing a multi-resolution autoregressive Huber loss in Time-MoE [timemoe]. As these models are pretrained for time series forecasting, their loss functions are directly reused during LoRA during fine-tuning on downstream datasets. Consequently, LoRA inherits the same network architecture, training objectives, and potential drawbacks of the original TSFM. Take Time-MOE [timemoe] as an example. Its multi-resolution loss function is defined as
@@
where $\mathcal{L}_{\text{ar}}$ denotes the Huber loss [huberloss], $P$ is the number of different prediction resolutions, $\alpha$ is a scaling factor, and $\mathcal{L}_{\text{aux}}$ is an auxiliary loss for routing in the mixture-of-experts (MoE) layers [dai2024deepseekmoe, riquelme2021scaling]. When fine-tuning Time-MoE on a downstream dataset with LoRA, its training objective must continue to balance the Huber loss and auxiliary MoE loss. Furthermore, the pretraining objectives of many TSFMs are often designed to learn generalized patterns through coarse predictions [chronos]. While beneficial for generalization across multiple domains, this can be a significant drawback for financial applications, where subtle predictive signals are crucial for accurate forecasting under changing market regimes.
@@ 3.4 Network Architecture
RefineBridge trains a context encoder and a 1D U-Net denoiser in an end-to-end fashion. For the context encoder, we employ the DLinear [dlinear] model to decompose the context into its trend and seasonal components. The encoder learns the latent representation of the context window, $\mathbf{z}_{c}=\mathcal{E}(\mathbf{c})\in\mathbb{R}^{H\times d}$.
@@ 4.1 Datasets
We evaluated RefineBridge on three distinct financial assets, the S&P 500 index (equity market), WTI crude oil (commodities), and the EUR/USD exchange rate (forex). We sourced 20 years of daily data (2005-01-01 to 2025-01-01), yielding approximately 5,040 trading days per asset. For each asset, the data were partitioned temporally into train/validation/test (80%/10%/10%) splits. Training samples were generated using a sliding window of step size 1. Each sample consisted of a triplet $\{\mathbf{x}_{T},\mathbf{c},\mathbf{x}_{0}\}$. We adopted a context window length of $C=21$ trading days for short forecasting horizons ($H\in\{5,10\}$) and $C=252$ trading days for longer forecasting horizons ($H\in\{21,63,126\}$). Given that financial time series exhibit heavy tails and frequent outliers which violate Gaussian assumptions, we followed [li2024master] and applied median–MAD robust normalization for all time series.
@@ 4.2 Experimental Settings
Implementation Details. All models were trained with the AdamW optimizer ($lr=10^{-3}$, $\epsilon=10^{-8}$) and a batch size of 512.
For short prediction horizons ($H\leq 10$), a noise schedule of $\beta_{0}=0.01,\beta_{1}=50$ was used for rapid refinement, while for long horizons ($H\geq$ 21) we used $\beta_{0}=0.0001,\beta_{1}=0.02$ for finer refinements. Notably, each RefineBridge model consisted of only 2.32.3M parameters, which was less than the number of parameters used in the LoRA adapters (7.6M in Chronos, 2.4M in Moirai, and 2.7M in Time-MoE).
@@ 4.3 Results
Remark 1. A distinguishing feature of RefineBridge is that, unlike in other Schrödinger Bridge models [bridgetts, wang2025framebridge, li2025bridge] where the prior may be directly part of or very close to the ground truth target, RefineBridge refines less informative or even false (low-quality) priors from TSFMs and still manages to improve the forecasting performance. This is shown in Fig. 1 and Fig. 2, where RefineBridge successfully corrects a prior trending in a different direction from the ground-truth. Such performance demonstrates the robustness of bridge models and lays a solid foundation for utilising them as a post-processing module of TSFMs.

### v2

@@ 1 Introduction
Inspired by the scaling laws observed in foundation models for natural language processing [8] and computer vision [7], Time Series Foundation Models (TSFM) [1, 30, 25, 23, 20] have extended this scaling paradigm to time series forecasting, with models containing millions to billions of parameters. By scaling the transformer architecture [27] and pre-training on diverse time series datasets, TSFMs can learn universal temporal representations and have achieved state-of-the-art performance across a wide range of forecasting tasks.
@@
Despite the presence of financial time series in the training data of TSFMs, their prediction remains particularly challenging due to domain-specific statistical properties [12]. In particular, financial data often exhibit near-zero autocorrelation, strong non-stationarity, and heteroskedasticity which shifts across market regimes. As a result, TSFMs tend to default to mean-reverting forecasts and output the average of the context window irrespective of market conditions. This leads to predictions that miss higher-frequency variations, which frequently carry tradable signals.
@@
Fig. 1: RefineBridge learns an optimal transport map from TSFM predictions to the ground truth via Schrödinger Bridge. At inference time, RefineBridge progressively refines the TSFM forecasts.
@@
Low-Rank Adaptation (LoRA) [11] has emerged as a popular parameter-efficient method for adapting TSFMs to downstream data domains [10]. LoRA-based adapters introduce low-rank update matrices with fewer trainable parameters than the original foundation model, enabling parameter-efficient fine-tuning. However, when applied to financial data, LoRA is prone to overfitting and can even degrade TSFM performance (see Table 1). Specifically, LoRA typically adopts the same network architecture and objective function as the original model, rather than introducing a complementary objective to enhance predictions, inherently limiting its performance.
@@
To address these issues, we propose RefineBridge, a novel paradigm for improving the performance of TSFMs on financial time series forecasting. In contrast to LoRA, which functions as a fine-tuning technique, RefineBridge operates as an independent post-processing module. Specifically, it treats TSFM forecasts as the prior distribution and iteratively refines them through Schrödinger Bridge optimal transport [4, 28, 17, 32], conditioned on the context window. As a probabilistic generative model, RefineBridge introduces a complementary training objective which captures the optimal stochastic transport maps from the prior to the target distribution. By using the coarse predictions by TSFMs as a prior, RefineBridge learns to model fine-grained details, which are crucial in financial time series forecasting. During the sampling stage, RefineBridge generates a refinement trajectory which aligns TSFM forecasts with the unique statistical properties of downstream financial data, all while keeping the original TSFM parameters unchanged, as illustrated in Fig. 1. The main contributions of this work are as follows.
@@ 2 Related Work
Time Series Foundation Models. Recent advancements in deep learning have led to the development of various TSFMs [1, 30, 25, 23, 20], with Chronos [1], Moirai [30], and Time-MoE [25] representing three distinct architectures with different training objective functions. Chronos [1] adapts the T5 [22] architecture by discretizing continuous time series into tokens and treating forecasting as a sequence-to-sequence task, where future tokens are generated auto-regressively. Moirai [30] instead uses mixture distributions and a masked encoder architecture with continuous embeddings. At a larger scale, Time-MoE [25] is the first TSFM to reach the billion-parameter level through a sparse mixture-of-experts architecture.
Low-rank Adaptation. Full fine-tuning of TSFMs for specific downstream datasets is computationally expensive due to their large parameter count. To address this, LoRA-based adapters [11, 15, 9] introduce low-rank weight update matrices to greatly reduce the number of trainable parameters during fine-tuning. However, in practice, LoRA struggles when applied to financial data, as shown in Table 1.
@@ 3 RefineBridge
Unlike diffusion models which map noise to data, Schrödinger Bridge learns data-to-data transport maps by computing the most probable stochastic evolution between two probability distributions under a reference diffusion process [3, 2]. This makes SB well-suited for refining TSFM predictions, which can be viewed as neural priors learned by the internal representation of TSFMs, as illustrated in Fig. 1. Even though TSFMs often struggle with financial time series, their forecasts still contain a coarse estimate (the neural prior) of the prediction target. As highlighted by [28], compared to noise-to-data diffusion models, SB-based models can capture the target distribution from a clean prior, making it possible to build post-processing refinement models for pretrained foundation models.
@@ 3.1 Bridging TSFM Predictions and Ground Truth
Following prior work [4, 28, 16], we adopt a tractable solution for paired data $(\mathbf{x}_{0},\mathbf{x}_{T})$. We set the reference stochastic differential equation (SDE) to have zero drift ($f(t)=0$) and a time-varying diffusion defined as $d\mathbf{x}_{t}=g(t)d\mathbf{w}_{t}$, where $g^{2}(t)=\beta_{0}+t(\beta_{1}-\beta_{0})$. Here, $f:\mathbb{R}^{H}\times[0,T]\to\mathbb{R}^{H}$ is the vector-valued drift term, $g:[0,T]\to\mathbb{R}$ is scalar-valued diffusion coefficient, and $\mathbf{w}_{t}\in\mathbb{R}^{H}$ denotes a standard Wiener process. This SDE defines the continuous-time diffusion process upon which score-based generative models (SGMs) [26, 5, 19, 21, 14, 13] are built.
In the general SB framework with linear drift, the scaling coefficients are given by $\alpha_{t}=e^{\int_{0}^{t}f(\tau)d\tau}$ and $\bar{\alpha}_{t}=e^{-\int_{t}^{1}f(\tau)d\tau}$. We employ the Bridge-gmax scheduler proposed in [4], to simplify these to $\alpha_{t}=\bar{\alpha}_{t}=1$ under $f(t)=0$. The corresponding variance schedules are $\sigma_{t}^{2}=\int_{0}^{t}g^{2}(\tau)d\tau=\beta_{0}t+\frac{1}{2}(\beta_{1}-\beta_{0})t^{2}$ and $\bar{\sigma}_{t}^{2}=\sigma_{1}^{2}-\sigma_{t}^{2}.$ The resulting bridge marginal distribution is Gaussian
@@ 3.2 Complementary Training Objective
When LoRA [11] is applied to a TSFM for downstream adaptation, it typically retains the training objective of the base model, thereby potentially inheriting its limitations. This includes maximizing token-level log-likelihood in Chronos [1], optimizing a mixture distribution log-likelihood in Moirai [30], and minimizing a multi-resolution autoregressive Huber loss in Time-MoE [25]. As these models are pretrained for time series forecasting, their loss functions are directly reused during LoRA during fine-tuning on downstream datasets. Consequently, LoRA inherits the same network architecture, training objectives, and potential drawbacks of the original TSFM. Take Time-MOE [25] as an example. Its multi-resolution loss function is defined as
@@
where $\mathcal{L}_{\text{ar}}$ denotes the Huber loss [29], $P$ is the number of different prediction resolutions, $\alpha$ is a scaling factor, and $\mathcal{L}_{\text{aux}}$ is an auxiliary loss for routing in the mixture-of-experts (MoE) layers [6, 24]. When fine-tuning Time-MoE on a downstream dataset with LoRA, its training objective must continue to balance the Huber loss and auxiliary MoE loss. Furthermore, the pretraining objectives of many TSFMs are often designed to learn generalized patterns through coarse predictions [1]. While beneficial for generalization across multiple domains, this can be a significant drawback for financial applications, where subtle predictive signals are crucial for accurate forecasting under changing market regimes.
@@ 3.4 Network Architecture
RefineBridge trains a context encoder and a 1D U-Net denoiser in an end-to-end fashion. For the context encoder, we employ the DLinear [31] model to decompose the context into its trend and seasonal components. The encoder learns the latent representation of the context window, $\mathbf{z}_{c}=\mathcal{E}(\mathbf{c})\in\mathbb{R}^{H\times d}$.
@@ 4.1 Datasets
We evaluated RefineBridge on three distinct financial assets, the S&P 500 index (equity market), WTI crude oil (commodities), and the EUR/USD exchange rate (forex). We sourced 20 years of daily data (2005-01-01 to 2025-01-01), yielding approximately 5,040 trading days per asset. For each asset, the data were partitioned temporally into train/validation/test (80%/10%/10%) splits. Training samples were generated using a sliding window of step size 1. Each sample consisted of a triplet $\{\mathbf{x}_{T},\mathbf{c},\mathbf{x}_{0}\}$. We adopted a context window length of $C=21$ trading days for short forecasting horizons ($H\in\{5,10\}$) and $C=252$ trading days for longer forecasting horizons ($H\in\{21,63,126\}$). Given that financial time series exhibit heavy tails and frequent outliers which violate Gaussian assumptions, we followed [18] and applied median–MAD robust normalization for all time series.
@@ 4.2 Experimental Settings
Implementation Details. All models were trained with the AdamW optimizer ($lr=10^{-3}$, $\epsilon=10^{-8}$) and a batch size of 512. For short prediction horizons ($H\leq 10$), a noise schedule of $\beta_{0}=0.01,\beta_{1}=50$ was used for rapid refinement, while for long horizons ($H\geq$ 21) we used $\beta_{0}=0.0001,\beta_{1}=0.02$ for finer refinements. Notably, each RefineBridge model consisted of only 2.32.3M parameters, which was less than the number of parameters used in the LoRA adapters (7.6M in Chronos, 2.4M in Moirai, and 2.7M in Time-MoE).
@@ 4.3 Results
Remark 1. A distinguishing feature of RefineBridge is that, unlike in other Schrödinger Bridge models [4, 28, 16] where the prior may be directly part of or very close to the ground truth target, RefineBridge refines less informative or even false (low-quality) priors from TSFMs and still manages to improve the forecasting performance. This is shown in Fig. 1 and Fig. 2, where RefineBridge successfully corrects a prior trending in a different direction from the ground-truth. Such performance demonstrates the robustness of bridge models and lays a solid foundation for utilising them as a post-processing module of TSFMs.